OpenStudy (anonymous):

A pastry recipe calls for 3 cups of flour to make 8 servings. How many cups of flour are needed to make 6 servings?

OpenStudy (anonymous):

Divide 3 by 8 to see that each serving takes .375, or 3/8ths cus of flour multiply this by 6 and you'll get your answer. In fraction or decimals.

OpenStudy (anonymous):

cups* of flour

OpenStudy (texaschic101):

3/8 = x/6 (3 cups to 8 servings = x cups to 6 servings) cross multiply 8x = 18 x = 18/8 or 2 1/4 cups

OpenStudy (anonymous):

This problem is similar to the previous one. It actually looks like this: 3/8 = N/6

OpenStudy (anonymous):

THANK YOU GOD I GOT A HUNDRED BECAUSE OF U GUYS i wanna give medal but i dont know who to

OpenStudy (anonymous):

Lol, you're welcome.